OpenStudy (anonymous):

Use the functions to answer the question. f(x) = 4x – 5; g(x) = x + 3 Which is equal to (g ◦ f)(x)?

OpenStudy (aum):

\[(g \circ f)(x) = g(f(x)) \\ g(x) = x + 3 \\ g(f(x)) = f(x) + 3 \\ f(x) = 4x - 5 \\ g(f(x)) = 4x - 5 + 3 = 4x - 2 \]

OpenStudy (anonymous):

thanks for answer :-D

OpenStudy (aum):

you are welcome.